​ For the function f defined by f(n) =n2+ 1/n+ 1 for n∈N, show that f(n)∈Θ(n). Use
1
Expert's answer
2020-10-14T18:24:16-0400

"f(n)=\dfrac{n^2+1}{n+1}\leq\dfrac{n^2+n}{n+1}=n, n\in \N"

"\N=\{1,2,...\}"

By the definition "f(n)" is "O(n)" with "C=1" and "k=1."



"\dfrac{n^2+1}{n+1}-\dfrac{n}{2}=\dfrac{2n^2+2-n^2-n}{2(n+1)}="

"=\dfrac{n^2-n+\dfrac{1}{4}+\dfrac{7}{4}}{2(n+1)}=\dfrac{(n-\dfrac{1}{2})^2+\dfrac{7}{4}}{2(n+1)}>0,n\in\N"


"|f(n|)=|\dfrac{n^2+1}{n+1}|>\dfrac{1}{2}

|n|, n\in\N"

Then "f(n)" is "\Omega(n)."

"\dfrac{1}{2}|n|<|f(n)|\leq|n|, n\in\N"

By the definition this means that "f(n)=\dfrac{n^2+1}{n+1}" is "\Theta(n), n\in\N."